OpenStudy (anonymous):

The data in the table relates the total cost y to have a party with x guests at a restaurant. In the data table, the x-column represents the number of guests and the y-column represents the total cost in dollars. When x is ten, y equals two hundred. When x is fifteen, y equals three hundred. When x is twenty-five, y equals five hundred. When x is thirty, y equals six hundred. Based on the information in the table, what is the function rule? y equals x plus one hundred ninety x equals y plus one hundred ninety y equals twenty x x equals twenty y

jigglypuff314 (jigglypuff314):

y = mx + b where m = slope and b = y-intercept first find the slope between two points (10, 200) and (15, 300) \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \]when given points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

Plug in the x given into each of the answer choices. You can eliminate b and d because those arent functions. So plug in 10 for x and 200 for y in both equations. 200=10+190, 200=20(10) those both work so plug in the next situation. What do you get?

OpenStudy (anonymous):

200=200?

OpenStudy (anonymous):

Yes but they both equal 200 so you can't pick answer yet because both of them wont necessarily work for the function. So now you plug in 15 for x and 300 for y into both functions what do you get?